OpenStudy (anonymous):

Express 0.40 x 0.50 as a fraction:

OpenStudy (anonymous):

1/5

jimthompson5910 (jim_thompson5910):

0.40 x 0.50 = 0.2 = 20/100 = 1/5

myininaya (myininaya):

\[0.4=4/10; 0.5=5/10 => \frac{4}{10}\cdot \frac{5}{10}=\frac{20}{100}=\frac{2 \cdot 10}{10 \cdot 10}=\frac{2}{10}=\frac{2}{2 \cdot 5}=\frac{1}{5}\]

myininaya (myininaya):

and we could had reduced those fractions before multiplying

myininaya (myininaya):

\[\frac{4}{10}=\frac{2 \cdot 2}{2 \cdot 5}=\frac{2}{5}; \frac{5}{10}=\frac{5}{2 \cdot 5}=\frac{1}{2}\] \[\frac{4}{10} \cdot \frac{5}{10}=\frac{2}{5} \cdot \frac{1}{2}=\frac{1}{5}\]

OpenStudy (anonymous):

thank u all i have a lot of them and not a clue of any of them